# Hyperbolic recurrent neural network as the first type of non-Euclidean neural quantum state ansatz

### v3

@@ Abstract
In this work, we introduce the first type of non-Euclidean neural quantum state (NQS) ansatz, in the form of the hyperbolic GRU (a variant of recurrent neural networks (RNNs)), to be used in the Variational Monte Carlo method of approximating the ground state wavefunction for quantum many-body systems. In particular, we examine the performances of NQS ansatzes constructed from both conventional or Euclidean RNN/GRU and from hyperbolic GRU in the prototypical settings of the one- and two-dimensional transverse field Ising models (TFIM) and the one-dimensional Heisenberg $J_{1}J_{2}$ and $J_{1}J_{2}J_{3}$ systems with varying values of $(J_{1},J_{2})$ and $(J_{1},J_{2},J_{3})$. By virtue of the fact that, for all of the experiments performed in this work, hyperbolic GRU can yield performances comparable to or better than Euclidean RNNs, which have been extensively studied in these settings in the literature, our work is a proof-of-concept for the viability of hyperbolic GRU as the first type of non-Euclidean NQS ansatz for quantum many-body systems. Furthermore, in settings where the Hamiltonian displays a clear hierarchical interaction structure, such as the 1D Heisenberg $J_{1}J_{2}$ & $J_{1}J_{2}J_{3}$ systems with the first, second and even third nearest neighbor interactions, our results show that hyperbolic GRU definitively outperforms its Euclidean version in almost all instances. Additionally, in the 2D TFIM setting, 1D hyperbolic GRU NQS also outperforms 1D Euclidean GRU NQS, due to the presence of a hierarchical interaction structure comprising the first and $N^{th}$ nearest neighbor interactions, introduced by the rearrangement of the 1D spin chain to mimic the 2D lattice. The fact that these results are reminiscent of the established ones from natural language processing where hyperbolic RNNs almost always outperform Euclidean RNNs when the training data exhibit a tree-like or hierarchical structure leads us to hypothesize that hyperbolic GRU NQS ansatz would likely outperform its Euclidean counterpart in quantum spin systems that involve different degrees of nearest neighbor interactions. Finally, with this work, we hope to initiate future studies of other types of non-Euclidean neural quantum states beyond hyperbolic GRU.
@@ 1 Introduction
With the advance of deep learning in physics, neural networks have been established as efficient variational trial wavefunctions (variational ansatzes) to estimate the ground states of quantum systems starting with the work of [1] that introduced restricted Boltzmann machines (RBM) as viable candidates for the trial wavefunctions in the spin model settings of 1D Ising model, 1D and 2D Heisenberg models. Since then, in addition to RBM being further employed as variational ansatz in various quantum systems [4], [13], many other types of neural networks have also been constructed and shown to be viable trial/variational wavefunctions in a variety of different problem settings in quantum and condensed matter physics ranging from spin systems [1] - [12], fermion systems [22]-[25], itinerant boson systems [20], [21] to matrix models [14], [40] in the recent literature. Some notable examples of these include ANN-based ansatzes [2], [3], CNN-based ansatzes [6], [7], [8], RNN-based ansatzes [9], [10], [11], [12], transformer-based ansatzes [15], [16] [17], [18], [19] and even generative networks like autoregressive [5] and normalizing-flow-based ansatzes [14], [40]. Collectively, this class of trial wavefunctions based on neural network models are known as neural quantum states (NQS). Very recently, a type of foundation NQS, based on the transformer architecture and introduced in the work [26], is the first type of versatile NQS capable of processing multimodal inputs comprised of spin configurations and Hamiltonian couplings, and generalizing beyond those Hamiltonians encountered during its training phase. Beyond the task of approximating the ground state, extensions of NQS for approximating the excited states are also an active direction of research [27]. A recent review of NQS, including the different neural network architectures used and their many applications in quantum many-body physics can be found in [28] and references therein.
@@
In this work, we present the first type of non-Euclidean NQS based on hyperbolic recurrent neural networks (RNNs). This type of hyperbolic RNNs, whose underlying hyperbolic geometry is that of a Poincaré ball, was introduced in the work of [30] in the context of natural language processing for the tasks of sentence embeddings, textual entailment and noisy-prefix recognition. For these tasks, hyperbolic RNNs have been shown to outperform their Euclidean counterparts, especially in settings when there the training data possess an inherent hierarchical structure [30], [33], [35]. Besides the Poincaré model, another variant of hyperbolic RNNs based on the Lorentz model [31] was proposed and deployed to carry out the NLP task of learning hierarchies in unstructured similarity scores. Besides hyperbolic RNNs, other types of hyperbolic networks, such as hyperbolic CNNs [36], hyperbolic attention networks [34], hyperbolic transformers [32], hyerbolic normalizing flows [33], also have been introduced in quick succession to tackle different NLP-related tasks in deep learning. In [32], the authors demonstrated the superior performances of Lorentz-based hyperbolic networks, including hyperbolic transformers, over their Euclidean versions, in the tasks of machine translation and knowledge graph completion. Regardless of the underlying hyperbolic model, i.e. either Poincaré or Lorentz, hyperbolic neural networks have consistently demonstrated their ability to better learn from and model data with latent hierarchies [30]-[35]. Beyond hyperbolic geometry, the possibility of constructing and deploying other neural networks based on different types of non-Euclidean geometries in deep learning has opened up new and exciting research directions for the relatively new field of non-Euclidean machine learning, whose examples include mixed-curvature variational autoencoders [37], constant curvature graph convolution neural networks [38] among others.
@@
Here, in the context of quantum many-body physics, we will reconfigure the hyperbolic GRU (Gated Recurrent Unit) as defined in [30] to be used as the first type of non-Euclidean NQS ansatz in the prototypical settings of the 1D and 2D transverse field Ising model (TFIM) as well as the 1D Heisenberg $J_{1}J_{2}$ and $J_{1}J_{2}J_{3}$ models. These models are the most basic testing grounds for a variety of experiments involving neural quantum states as demonstrated in the earliest works [1], [2] on this topic. With the aim of determining the viability and competitiveness of the hyperbolic GRU as an NQS ansatz, in each of these settings (Ising and Heisenberg models), the performance of the hyperbolic GRU ansatz will be benchmarked firstly against the DMRG (Density Matrix Renormalization Group) results and secondly against that of the conventional or Euclidean RNN/GRU ansatz introduced in the work of [9]. Our obtained results not only demonstrate this feasibility but also potentially open up the possibility of better performances with hyperbolic GRU in quantum settings that exhibit hierarchical structure. Although more work are needed to conclusively confirm this observation in quantum systems beyond those studied in this work, given these first promising results, we hope to highlight the potential of non-Euclidean NQS in quantum many-body physics.
@@
The Python codes used in this work to construct the hyperbolic GRU (using Tensorflow) [^2] For the construction of the hyperbolic GRU neural network, we adapted the Tensorflow-v1 (TF1) codes of [30] written for hyperbolic RNNs in the context of natural language processing tasks, which can be found at https://github.com/dalab/hyperbolic_nn, to the Tensorflow version 2 (TF2) used in this work and to run all VMC experiments [^3] For the VMC experiments as well as the construction of general RNN-based NQS ansatzes, we adapted the Tensorflow v1 (TF1) codes of [9] found at https://github.com/mhibatallah/RNNWavefunctions/ to Tensorflow v2 (TF2) with the addition of the hyperbolic GRU ansatz. as well as all the trained neural networks can be found at this Github repository: https://github.com/lorrespz/nqs_hyperbolic_rnn.
@@ 3.3 From Euclidean GRU to Hyperbolic GRU
    $\vec{\tilde{h}}_{i}=f^{\otimes_{c}}\left[(W\,\text{diag}(r_{t}))\otimes_{c}\vec{h}_{i-1}\oplus(U\otimes_{c}\vec{x}_{i})\frac{}{}\oplus_{c}\vec{b}_{h}\right]$
@@
$\vec{\tilde{h}}_{i}$ $=$ $f^{\otimes_{c}}\left[(W\,\text{diag}(r_{t}))\otimes_{c}\vec{h}_{i-1}\oplus(U\otimes_{c}\vec{x}_{i})\frac{}{}\oplus_{c}\vec{b}_{h}\right]\,$, (32)
@@ 4.1 Real NQS wavefunction
    and the Softmax function999which converts a vector of length K into a probability distribution of $K$ possible outcomes is
@@ 5 VMC with hyperbolic RNN-based NQS
  • In a typical training process, the weights of the neural network model under training are constantly overwritten after each training epoch, so the final model is the one saved at the last epoch in the training. Theoretically, at convergence when the model reaches a constant value without any further changes, the weights of the model stop changing and the weights at the last epoch are the final weights of the model. In practice, what is observed is that even at convergence when the model has reached the constant energy level, there are still minor fluctuations within a very small range of this energy, espcially with this kind of training involving the generation of samples. To address this issue, in our training process, the saving of the model weights at any particular training epoch is conditioned on the fact that the model has actually improved (the criteria for judging the improvement, programmed in the training loop, include not only a lower mean energy but also a lower variance, subject to a certain variance tolerance threshold) compared to the previous training epoch. As such, when convergence is reached in a practical sense (in the sense that the mean energy has reached a constant level albeit with fluctuations within a very narrow range), our last saved model correponds to the best possible result, i.e. a model that has already converged with the lowest energy and the variance smaller than the tolerance threshold. This means that our final model might not necessarily be the model at the last epoch in the training, but the model saved at some earlier epoch near the end of the training [^12] We note that the practice of saving the best model only after a practical convergence is reached is commonly used in all types of neural network training not restricted to the present context of NQS. In Keras (keras.io/api/callbacks/model_checkpoint/), one can choose to save the best model only (e.g. one that has the lowest validation loss or one that fulfills a certain set of custom defined criteria) as the training progresses using a function built-in in ModelCheckpoint callback by selecting the option save_best_only = True. . In other words, the training loop might continue until the specified number of iterations is reached but unless a better result is obtained, the final model is the model saved at the epoch where the best result was obtained [^13] Our training results are perfectly reproducible, as long as the same random seed parameter is chosen. . For example, with the total number of epochs set at 450, if the convergence is reached after 350 epochs (the energy no longer changes noticeably but varies within a very narrow range of order 0.01 to 0.1), and the best model (with the lowest energy and the variance below the threshold) is the one at epoch 375, training will still carry on until epoch 450 but no new model is saved after epoch 375 unless there is an improved model at the $i^{th}$ epoch where $375<i<450$.
@@ 5.1.1 Experiment details
The number of parameters of the Euclidean RNN ansatz is 2752, while the number of parameters of the Euclidean and hyperbolic GRU-based ansatz is 8052. For all three types of ansatzes at all $N$, convergence is reached very quickly at around 100 epochs with the total number of training steps set at 120 epochs. Even with this very small number of epochs, for the hyperbolic GRU to complete the training loop, it takes more than 1.5 hours for $N=20$ and up to 20 hours for $N=100$ on an average laptop141414a average laptop at the time of writing is one with around 8-16Gb with no GPU. On the other hand, Euclidean RNN and GRU take around 0.1 hours for $N=20$ and around 1.5 hours for $N=100$ on the same hardware.
@@ 5.2.2 Results
This result is very interesting since it might be indicative of the role played by the hyperbolic geometry of the NQS ansatz in Hamiltonian systems where there exist some sort of hierarchical structure. In the context of natural language processing (NLP) when the hyperbolic RNN/GRU was introduced in the work of [30], the authors noted a definitive and clear outperformance of hyperbolic RNN/GRU over their Euclidean version when the underlying data exhibit a hierarchical, tree-like structure. In the present 2D TFIM setting, there exists a hierarchy in the interaction structure of 2D Hamiltonian when 1D NQS is used.
This is due to the artificial rearrangement (or folding) of the 1D spin chain samples generated by the 1D NQS to model the 2D data during the calculations of their local energies (see Fig.10 in Appendix A). This process of rearrangement takes into account interactions between neighbors at site $i$ and site $i+N$, which were originally faraway in the 1D horizontal chain but are brought close together to form the vertical dimension in the 2D square lattice. Concretely, let us consider a chain of length $N^{2}=(N-1)N+N$ spins (see Subfigure (a) of Fig.10) folded into a square lattice of size ($N,N$) in Subfigure (b) of Fig.10. In 1D when only the next nearest neighbor interactions are counted, site $i$ and site $i+N$ have no interaction, but when arranged in a 2D square lattice, these sites are each other next nearest neighbor vertically, which means their interactions are counted in the local energy calculations. Together with the original 1D next nearest neighbor interactions $\langle i,i+1\rangle$, these additional interactions $\langle i,i+N\rangle$ actually form a hierarchy of interactions [^16] Note that in the folding process of the 1D spin chain to model the 2D data, alongside the ‘gained’ (in the sense that these are included in the Hamiltonian calculations due to the 2D geometry) interaction $\langle i,i+N\rangle$, we also ‘lose’ some next neareast neighbor interactions $\langle i,i+1\rangle$ when $i=kN$, with $k\in\mathbb{N}$, i.e. when $i$ is multiple of $N$., and it is this hierarchy that might have played a part in the observation of 1D hyperbolic GRU outperforming 1D Euclidean GRU in 2D TFIM setting.

### v4

@@ Abstract
In this work, we introduce the first type of non-Euclidean neural quantum state (NQS) ansatz, in the form of the hyperbolic GRU (a variant of recurrent neural networks (RNNs)), to be used in the Variational Monte Carlo method of approximating the ground state wavefunction for quantum many-body systems. In particular, we examine the performances of NQS ansatzes constructed from both conventional or Euclidean RNN/GRU and from hyperbolic GRU in the prototypical settings of the one- and two-dimensional transverse field Ising models (TFIM) and the one-dimensional Heisenberg $J_{1}J_{2}$ and $J_{1}J_{2}J_{3}$ systems with varying values of $(J_{1},J_{2})$ and $(J_{1},J_{2},J_{3})$. By virtue of the fact that, for all of the experiments performed in this work, hyperbolic GRU can yield performances comparable to or better than Euclidean RNNs, which have been extensively studied in these settings in the literature, our work is a proof-of-concept for the viability of hyperbolic GRU as the first type of non-Euclidean NQS ansatz for quantum many-body systems. Furthermore, in settings where the Hamiltonian displays a clear hierarchical interaction structure, such as the 1D Heisenberg $J_{1}J_{2}$ & $J_{1}J_{2}J_{3}$ systems with the first, second and even third nearest neighbor interactions, our results show that hyperbolic GRU definitively outperforms its Euclidean version in almost all instances. Additionally, in the 2D TFIM setting, 1D hyperbolic GRU NQS also outperforms 1D Euclidean GRU NQS, due to the presence of a hierarchical interaction structure comprising the first and $N^{th}$ nearest neighbor interactions, introduced by the rearrangement of the 1D spin chain to mimic the 2D $N\times N$ lattice. The fact that these results are reminiscent of the established ones from natural language processing where hyperbolic RNNs almost always outperform Euclidean RNNs when the training data exhibit a tree-like or hierarchical structure leads us to hypothesize that hyperbolic GRU NQS ansatz would likely outperform its Euclidean counterpart in quantum spin systems that involve different degrees of nearest neighbor interactions. Finally, with this work, we hope to initiate future studies of other types of non-Euclidean neural quantum states beyond hyperbolic GRU.
@@ 1 Introduction
With the advance of deep learning in physics, neural networks have been established as efficient variational trial wavefunctions (variational ansatzes) to estimate the ground states of quantum systems starting with the work of [1] that introduced restricted Boltzmann machines (RBM) as viable candidates for the trial wavefunctions in the spin model settings of 1D Ising model, 1D and 2D Heisenberg models. Since then, in addition to RBM being further employed as variational ansatz in various quantum systems [4], [13], many other types of neural networks have also been constructed and shown to be viable trial/variational wavefunctions in a variety of different problem settings in quantum and condensed matter physics ranging from spin systems [1] - [12], fermion systems [22]-[25], itinerant boson systems [20], [21] to matrix models [14], [46] in the recent literature. Some notable examples of these include ANN-based ansatzes [2], [3], CNN-based ansatzes [^2] ANN stands for Artifical Neural Network, while CNN stands for Convolutional Neural Network. [6], [7], [8], RNN-based ansatzes [9], [10], [11], [12], transformer-based ansatzes [15], [16] [17], [18], [19] and even generative networks like autoregressive [5] and normalizing-flow-based ansatzes [14], [46]. Collectively, this class of trial wavefunctions based on neural network models are known as neural quantum states (NQS). Very recently, a type of foundation NQS, based on the transformer architecture and introduced in the work [26], is the first type of versatile NQS capable of processing multimodal inputs comprised of spin configurations and Hamiltonian couplings, and generalizing beyond those Hamiltonians encountered during its training phase. Beyond the task of approximating the ground state, extensions of NQS for approximating the excited states are also an active direction of research [27]. A recent review of NQS, including the different neural network architectures used and their many applications in quantum many-body physics can be found in [28] and references therein.
@@
In this work, we present the first type of non-Euclidean NQS based on hyperbolic recurrent neural networks (RNNs). This type of hyperbolic RNNs, whose underlying hyperbolic geometry is that of a Poincaré ball, was introduced in the work of [30] in the context of natural language processing for the tasks of sentence embeddings, textual entailment and noisy-prefix recognition. For these tasks, hyperbolic RNNs have been shown to outperform their Euclidean counterparts, especially in settings when there the training data possess an inherent hierarchical structure [30], [34], [32], [36]. Besides the Poincaré model, another variant of hyperbolic RNNs based on the Lorentz model [32] was proposed and deployed to carry out the NLP task of learning hierarchies in unstructured similarity scores. Besides hyperbolic RNNs, other types of hyperbolic networks, such as hyperbolic CNNs [37], hyperbolic attention networks [35], hyperbolic transformers [33], hyperbolic normalizing flows [34], also have been introduced in quick succession to tackle different NLP-related tasks in deep learning. In [33], the authors demonstrated the superior performances of Lorentz-based hyperbolic networks, including hyperbolic transformers, over their Euclidean versions, in the tasks of machine translation and knowledge graph completion. Regardless of the underlying hyperbolic model, i.e. either Poincaré or Lorentz, hyperbolic neural networks have consistently demonstrated their ability to better learn from and model data with latent hierarchies [30]-[36]. Beyond hyperbolic geometry, the possibility of constructing and deploying other neural networks based on different types of non-Euclidean geometries in deep learning has opened up new and exciting research directions for the relatively new field of non-Euclidean machine learning, whose examples include mixed-curvature variational autoencoders [38], constant curvature graph convolution neural networks [39] among others.
In the aforementioned tasks involving hierarchical, tree-like data structures, the obvious outperformances delivered by hyperbolic neural networks when benchmarked against their Euclidean counterparts can be largely attributed to the hyperbolic geometry underlying the construction of these networks. It was known from the works of [40], [41], [42], [43] and [44] that hyperbolic space has the capacity to embed or represent tree structures with very low distortion and arbitrarily small errors. This is because in $D$-dimensional hyperbolic space $\mathbb{H}_{D}$, the volume grows exponentially with respect to the radius, while in $D$-dimensional Euclidean space $\mathbb{E}_{D}$, the volume only grows polynomially [^3] To be concrete, let’s consider the case of 2D hyperbolic space $\mathbb{H}_{2}$ and 2D Euclidean space $\mathbb{E}_{2}$. In $\mathbb{H}_{2}$ with the negative constant curvature $\zeta$ where $\zeta>0$, the length of the circle and the area of the disk of radius $R$ are $2\pi\sinh(\zeta R)$ and $2\pi\,(\cosh\zeta R-1)$, respectively. Both of these quantities grow as $e^{\zeta R}$, i.e. exponentially with respect to $R$. In Euclidean space $\mathbb{E}_{2}$, the circle length and disk area of radius $R$ are $2\pi R$ and $\pi R^{2}$, respectively. These only grow polynomially with respect to $R$. . Due to the exponential expansion of space in hyperbolic geometry, there exists a direct link or map between a $b$-ary tree (a tree with branching factor $b$) and hyperbolic space, with the former informally thought of as a discrete version of the latter [^4] This is the case because the number of nodes at a distance $R$ away from the root of a $b$-ary tree is $(b+1)b^{R}$, which grows as $b^{R}$ as $R$ increases, similar to the circle length in hyperbolic space. As such, hyperbolic spaces allow for almost isometric embeddings of tree structures [^5] In particular, the work [41] showed that tree structures can be embedded without distortion in 2D hyperbolic space, a Poincaré disk, while the work [44] extended this embedding to $D$-dimensional hyperbolic space., which require an exponential amount of space for branching, while Euclidean spaces (of any dimensions) do not [40], [41], [43], [44].
@@
Here, in the context of quantum many-body physics, we will reconfigure the hyperbolic GRU (Gated Recurrent Unit) as defined in [30] to be used as the first type of non-Euclidean NQS ansatz in the prototypical settings of the 1D and 2D transverse field Ising model (TFIM) as well as the 1D Heisenberg $J_{1}J_{2}$ and $J_{1}J_{2}J_{3}$ models. These models are the most basic testing grounds for a variety of experiments involving neural quantum states as demonstrated in the earliest works [1], [2] on this topic. With the aim of determining the viability and competitiveness of the hyperbolic GRU as an NQS ansatz, in each of these settings (Ising and Heisenberg models), the performance of the hyperbolic GRU ansatz will be benchmarked firstly against the DMRG (Density Matrix Renormalization Group) results and secondly against that of the conventional or Euclidean RNN/GRU ansatz introduced in the work of [9]. Our obtained results not only demonstrate this feasibility but also potentially open up the possibility of better performances with hyperbolic GRU in quantum settings that exhibit hierarchical structures. Although more work are needed to conclusively confirm this observation in quantum systems beyond those studied in this work, given these first promising results, we hope to highlight the potential of non-Euclidean NQS in quantum many-body physics.
@@
The Python codes used in this work to construct the hyperbolic GRU (using Tensorflow) [^6] For the construction of the hyperbolic GRU neural network, we adapted the Tensorflow-v1 (TF1) codes of [30] written for hyperbolic RNNs in the context of natural language processing tasks, which can be found at https://github.com/dalab/hyperbolic_nn, to the Tensorflow version 2 (TF2) used in this work and to run all VMC experiments [^7] For the VMC experiments as well as the construction of general RNN-based NQS ansatzes, we adapted the Tensorflow v1 (TF1) codes of [9] found at https://github.com/mhibatallah/RNNWavefunctions/ to Tensorflow v2 (TF2) with the addition of the hyperbolic GRU ansatz.
as well as all the trained neural networks can be found at this Github repository: https://github.com/lorrespz/nqs_hyperbolic_rnn.
@@ 3.3 From Euclidean GRU to Hyperbolic GRU
    $\vec{\tilde{h}}_{i}=f^{\otimes_{c}}\left[(W\,\text{diag}(r_{i}))\otimes_{c}\vec{h}_{i-1}\oplus(U\otimes_{c}\vec{x}_{i})\frac{}{}\oplus_{c}\vec{b}_{h}\right]$
@@
$\vec{\tilde{h}}_{i}$ $=$ $f^{\otimes_{c}}\left[(W\,\text{diag}(r_{i}))\otimes_{c}\vec{h}_{i-1}\oplus(U\otimes_{c}\vec{x}_{i})\frac{}{}\oplus_{c}\vec{b}_{h}\right]\,$, (32)
@@ 4.1 Real NQS wavefunction
    and the Softmax function131313which converts a vector of length K into a probability distribution of $K$ possible outcomes is
@@ 5 VMC with hyperbolic RNN-based NQS
  • In a typical training process, the weights of the neural network model under training are constantly overwritten after each training epoch, so the final model is the one saved at the last epoch in the training. Theoretically, at convergence when the model reaches a constant value without any further changes, the weights of the model stop changing and the weights at the last epoch are the final weights of the model. In practice, what is observed is that even at convergence when the model has reached the constant energy level, there are still minor fluctuations within a very small range of this energy, espcially with this kind of training involving the generation of samples. To address this issue, in our training process, the saving of the model weights at any particular training epoch is conditioned on the fact that the model has actually improved (the criteria for judging the improvement, programmed in the training loop, include not only a lower mean energy but also a lower variance, subject to a certain variance tolerance threshold) compared to the previous training epoch. As such, when convergence is reached in a practical sense (in the sense that the mean energy has reached a constant level albeit with fluctuations within a very narrow range), our last saved model correponds to the best possible result, i.e. a model that has already converged with the lowest energy and the variance smaller than the tolerance threshold. This means that our final model might not necessarily be the model at the last epoch in the training, but the model saved at some earlier epoch near the end of the training [^16] We note that the practice of saving the best model only after a practical convergence is reached is commonly used in all types of neural network training not restricted to the present context of NQS. In Keras (keras.io/api/callbacks/model_checkpoint/), one can choose to save the best model only (e.g. one that has the lowest validation loss or one that fulfills a certain set of custom defined criteria) as the training progresses using a function built-in in ModelCheckpoint callback by selecting the option save_best_only = True. . In other words, the training loop might continue until the specified number of iterations is reached but unless a better result is obtained, the final model is the model saved at the epoch where the best result was obtained [^17] Our training results are perfectly reproducible, as long as the same random seed parameter is chosen. . For example, with the total number of epochs set at 450, if the convergence is reached after 350 epochs (the energy no longer changes noticeably but varies within a very narrow range of order 0.01 to 0.1), and the best model (with the lowest energy and the variance below the tolerance threshold) is the one at epoch 375, training will still carry on until epoch 450 but no new model is saved after epoch 375 unless there is an improved model at the $i^{th}$ epoch where $375<i<450$.
@@ 5.1.1 Experiment details
The number of parameters of the Euclidean RNN ansatz is 2752, while the number of parameters of the Euclidean and hyperbolic GRU-based ansatz is 8052. For all three types of ansatzes at all $N$, convergence is reached very quickly at around 100 epochs with the total number of training steps set at 120 epochs. Even with this very small number of epochs, for the hyperbolic GRU to complete the training loop, it takes more than 1.5 hours for $N=20$ and up to 20 hours for $N=100$ on an average laptop [^18] An average laptop at the time of writing is one with around 8-16Gb of RAM with no GPU. On the other hand, Euclidean RNN and GRU take around 0.1 hours for $N=20$ and around 1.5 hours for $N=100$ on the same hardware.
@@ 5.2.2 Results
This result is very interesting since it might be indicative of the role played by the hyperbolic geometry of the NQS ansatz in Hamiltonian systems when there exist some sort of hierarchical structure. We recall that in the context of natural language processing (NLP) when the hyperbolic RNN/GRU was introduced in the work of [30], the authors noted a definitive and clear outperformance of hyperbolic RNN/GRU over their Euclidean version when the underlying data exhibit a hierarchical, tree-like structure. In the present 2D TFIM setting, there exists a hierarchy in the interaction structure of 2D Hamiltonian when 1D NQS is used. This is due to the artificial rearrangement (or folding) of the 1D spin chain samples generated by the 1D NQS to model the 2D data during the calculations of their local energies (see Fig.10 in Appendix A). This process of rearrangement takes into account interactions between neighbors at site $i$ and site $i+N$, which were originally faraway in the 1D horizontal chain but are brought close together to form the vertical dimension in the 2D square lattice.
Concretely, if we start with a two-dimensional $N\times N$ square lattice of spins whose sites are labeled as $(i^{\prime},j^{\prime})$ where $1\leq i^{\prime}\leq N$ and $1\leq j^{\prime}\leq N$, the 2D TFIM Hamiltonian of this system contains the nearest-neighbor interactions $\langle(i^{\prime},j^{\prime}),(i^{\prime},j^{\prime}+1)\rangle$ along the horizontal dimension as well as $\langle(i^{\prime},j^{\prime}),(i^{\prime}+1,j^{\prime})\rangle$ along the vertical dimension. When treated as a one-dimensional system, the 2D lattice can be thought of as a chain of length $N^{2}=(N-1)N+N$ spins (see Subfigure (a) of Fig.10) folded into a square lattice of size ($N,N$) (see Subfigure (b) of Fig.10). The mapping of the site location in going from the 2D square spin lattice to the 1D spin chain system is:
$\text{2D}:\,(i^{\prime},j^{\prime})\,\,\,\rightarrow\,\,\text{1D}:\,\left((i^{\prime}-1)N+j^{\prime}\right)\,.$ (53)
This leads to the original 2D nearest-neighbor interactions becoming the following 1D interactions:
$\text{2D}:\,\left\langle(i^{\prime},j^{\prime}),\frac{}{}(i^{\prime},j^{\prime}+1)\right\rangle\,\,$ $\rightarrow$ $\,\,\text{1D}:\,\left\langle(i^{\prime}-1)N+j^{\prime},\frac{}{}(i^{\prime}-1)N+j^{\prime}+1\right\rangle$
$\text{2D}:\,\left\langle(i^{\prime},j^{\prime}),\frac{}{}(i^{\prime}+1,j^{\prime})\right\rangle\,\,$ $\rightarrow$ $\,\,\text{1D}:\,\left\langle(i^{\prime}-1)N+j^{\prime},\frac{}{}i^{\prime}N+j^{\prime}\right\rangle$ (54)
Effectively, with a simple redefinition of the index $i=(i^{\prime}-1)N+j^{\prime}$, these 1D interactions are:
$\left\langle i,i+1\right\rangle,\hskip 14.22636pt\langle i,i+N\rangle\,.$
This means that in translating the 2D spin lattice to the 1D spin chain so that the 1D Euclidean and hyperbolic NQS ansatzes can be used (instead of the 2D Euclidean RNN ansatz), the original nearest-neighbor interactions along the horizontal and vertical dimensions of the 2D lattice become the nearest-neighbor $\langle i,i+1\rangle$ and the $N^{th}$-nearest neighbor $\langle i,i+N\rangle$ interactions in the 1D chain system. A hierarchy of interactions is thus formed when both the the nearest and $N^{th}$-nearest neighbor interactions are present, and it is this hierarchy that might have played a part in the observation of 1D hyperbolic GRU outperforming 1D Euclidean GRU in this 2D TFIM setting.